A shopper buys three oranges and five lemons for $10.26, while a second shopper buys four lemons and six oranges for $11.16. What is the price of each fruit?

Answer by hydromojo(20)

3O+5L = $10.26 ............#1
6O+4L = $11.16 ............#2
divide #2 by 2 throughout, we get
3O+2L = $5.58 ...........#3
subtract #2 from #1, you get the price of oranges as 82c
put this value in either equation, you get the value of lemons as $1.56